lella ------ loint Coordinates Label = 3/9-- - MOO-> olololo Yin BOX
calculate all support reactions and internal forces of each member by hand and show working.

10lb


= = = = = = = = = ه ه ه ه ه ه = - - = = = = هو 7 ==== = == = ه ه ه ه ه ه ه ه مه حه سه مه وه
Answer #1

2016 lb GI E 14 D I F T prolb 2-625. 5-25 L 10.5 21 sign Convention towards the point > compression (C) - ve away from the juJoint E 67 61 A 41.190 (Sv=0) FGE S?n41-19 = FEHSEN 41-19 FGE SER FGE= Fen fen= 0.00666) FEG + FEGC0341-19 = FEH, & FEHC0541-Joent G FG, G=0 FAGI = FGIE = 8.75(T) re Joint G 41-19° 41-199 FAG > 13.2885 (Evo) FAG Sim 48.81 = FEG Los 48.818 + FEGC0341-Joint Hi Iso FEM 213-305 - FODH, (T) / | Ful, o ont H Joint H 15-19 H 49.41 4-190 / 48.86 4 80 0-00 (Ev=0) FELSO 48.81 t ForSo assinned direction is wrong. (for a 15.186 (T) FH = 0.0094(T) 20 Toint c Joint a | Foga 2010 Now since the trusses symmetr